### v4

@@ 4 Damped Iterative Reweighted $\ell_{2}$ Algorithms
The ${\rm IRL}_{2}$ algorithm is an effective method to solve problem (P). In this method, we write $h(\cdot)=r\circ\sqrt{(\cdot)}$, so that $s(x_{i})=h(x^{2}_{i})$. Since $\sqrt{\cdot}$ is concave on $\mathbb{R}_{+}$, $\sqrt{\lambda y_{1}+(1-\lambda)y_{2}}\geq\lambda\sqrt{y_{1}}+(1-\lambda)\sqrt{y_{2}}$ for any $\lambda\in[0,1]$ and $y_{1},y_{2}\geq 0$. Since $r$ is increasing and concave, it follows that $h(\lambda y_{1}+(1-\lambda)y_{2})=r(\sqrt{\lambda y_{1}+(1-\lambda)y_{2}})\geq r(\lambda\sqrt{y_{1}}+(1-\lambda)\sqrt{y_{2}})\geq\lambda r(\sqrt{y_{1}})+(1-\lambda)r(\sqrt{y_{2}})=\lambda h(y_{1})+(1-\lambda)h(y_{2})$, meaning $h$ is also concave on $\mathbb{R}_{+}$.
At each iteration, after adding perturbation $\epsilon_{i}^{2}$ to each $h$ and linearize $h$ at $(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}$, the ${\rm IRL}_{2}$ algorithm solves the following weighted $\ell_{2}$ regularized subproblem (for brievty, we still use $S$ and $T$ for the maps),
$y^{k}=S^{x}(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2},$
@@
$w(x_{i},\epsilon_{i})=h^{\prime}(x_{i}^{2}+\epsilon_{i}^{2})=\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{2\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}.$
In particular,
$w(x_{i},0)=\begin{cases}\frac{r^{\prime}(|x_{i}|)}{2|x_{i}|}&\text{ if }x_{i}\neq 0\\ \infty&\text{ if }x_{i}=0.\end{cases}$
@@
$y^{k}_{i}=S_{i}^{x}(x,\epsilon):=\frac{1}{1+\frac{2\lambda}{\beta}w(x_{i}^{k},\epsilon_{i}^{k})}(x_{i}^{k}-\frac{1}{\beta}\nabla_{i}f(x^{k})).$ (33)
If $w(x_{i}^{k},\epsilon_{i}^{k})=\infty$, we simply set $S_{i}^{x}(x^{k},\epsilon^{k})=0$. The damped iteratively reweighted $\ell_{2}$ algorithm, hereinafter named DIRL2 algorithm, is presented in Algorithm 10.
1:Input $\beta>\alpha L_{f}/2$, $\alpha\in(0,1)$, $x_{0}$, and $\epsilon_{0}>0$.
@@
4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=h^{\prime}((x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2})$.
@@
6:     $y^{k}=\arg\min\limits_{y\in\mathbb{R}^{n}}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2}.$
@@
The fixed point iteration of DIRL2 is then given as $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})$ with $T=(1-\alpha)I+\alpha S$ and
$T(x,\epsilon):=\begin{bmatrix}(1-\alpha)x+\alpha S^{x}(x,\epsilon)\\ (1-\alpha(1-\mu))\epsilon\end{bmatrix}.$ (34)
We have the following result to connect the stationary point of (P) and the fixed point of $T$.
Proposition 16.
If $x^{*}$ is a stationary point of (P), then $x^{*}=S^{x}(x^{*},0)$ and $(x^{*},0)=T(x^{*},0)$. Conversely, for any $x^{*}$ satisfying $x^{*}=S^{x}(x^{*},0)$, if $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, then $x^{*}$ is a stationary point of (P).
Proof.
Suppose $x^{*}$ is a stationary point. It follows from (6) that for $i\in\mathcal{I}^{*}$, $x_{i}^{*}=x_{i}^{*}-\tfrac{1}{\beta}(\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*}))$ implying $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$. Therefore, $x_{i}^{*}=S_{i}(x^{*},0)$ since $w(x_{i}^{*},0)=\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}$. As for $i\in\mathcal{J}^{*}$, $w(0,0)=\infty$, implying that $0=S_{i}(x^{*},0)$. In summary, $x^{*}=S^{x}(x^{*},0)$.
Now consider $x^{*}=S^{x}(x^{*},0)$. If $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, we only have to verify (6) is satisfied by $x^{*}$. It follows from $x^{*}_{i}=S_{i}(x^{*},0),i\in\mathcal{I}^{*}$ that $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$, implying $\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*})=0$, completing the proof. ∎
The converse statement of Proposition 16 may not be true when $\mathcal{J}^{*}\neq\emptyset$ and $r^{\prime}(|0^{+}|)<\infty$. Since for any $x^{*}$ that satisfies $x_{i}^{*}=S_{i}(x^{*},0)$, $i\in\mathcal{I}^{*}$, it is naturally true that $w(0,0)=\infty$ and $S_{i}(x^{*},0)=0$. However, there is no guarantee that $x^{*}$ satisfies (7). In those cases, $S$ may have two types of fixed points — those are stationary points and others are not. This represents a stark difference from the DIRL1 algorithm, where all fixed points of the iteration map are stationary points.
@@ 4.1 Convergence analysis
Proposition 17.
Suppose Assumption 1-3 hold. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$. It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and the reduction satisfies
$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
Proof.
@@
$r(\sqrt{(x_{i}^{k+1})^{2}+(\epsilon_{i}^{k})^{2}})\leq r(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})+w(x_{i}^{k},\epsilon^{k}_{i})((x_{i}^{k+1})^{2}-(x_{i}^{k})^{2}).$
@@
$\Delta F(x^{k+1},\epsilon^{k+1})\geq\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}.$ (35)
For each subproblem, $y^{k}$ satisfies the optimal condition
$\nabla f(x^{k})+2\lambda W(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})=0.$ (36)
Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimal condition (27), the following holds
$\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$ (37)
$=$ $\ \nabla f(x^{k})^{T}(x^{k}-x^{k+1})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})((x_{i}^{k})^{2}-(x_{i}^{k+1})^{2})-\frac{\beta}{2}\|x^{k}-x^{k+1}\|^{2}$
$\geq$ $\ \nabla f(x^{k})^{T}(\alpha x^{k}-\alpha y^{k})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(\alpha(x_{i}^{k})^{2}-\alpha(y_{i}^{k})^{2})-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$\geq$ $\ \alpha[\nabla f(x^{k})^{T}(x^{k}-y^{k})+2\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}(x_{i}^{k}-y_{i}^{k})]-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ \alpha[\nabla f(y^{k})+2\lambda W(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})]^{T}(x^{k}-y^{k})+\frac{\alpha\beta(2-\alpha)}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2},$
where the first inequality is by Lemma A.2 and the second is by the convexity of $(\cdot)^{2}$. Combining (26) and (37), we get
$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
@@
$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
$(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, from the boundedness of $\{x^{k}\}$, there must exist $C>0$ such that (23) holds for any $k\in\mathbb{N}$. ∎
Theorem 18.
(Global convergence) Suppose Assumption 13 hold, and $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{2}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then any $x^{*}\in\Omega$ is first-order optimal for (P).
Proof.
Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. The optimality condition of the $k$th subproblem implies
$-\beta(y_{i}^{k}-x_{i}^{k})=\nabla_{i}f(x^{k})+2\lambda w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}=\nabla_{i}f(x^{k})+\frac{r^{\prime}(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})}{\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}}}y_{i}^{k}.$ (38)
As $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$, $y_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$, $\lim\limits_{k\in\mathcal{S}}\frac{y_{i}^{k}}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}=\text{sign}(x_{i}^{*})$ for $i\in\mathcal{I}^{*}$.
For $i\in\mathcal{J}^{*}$, there is nothing to prove since $|\nabla_{i}f(x^{*})|\leq r^{\prime}(|0^{+}|)$ naturally holds. In general, $x^{*}$ is stationary. ∎

### v5

@@ 4 Damped Iterative Reweighted $\ell_{2}$ Algorithms
The ${\rm IRL}_{2}$ algorithm is an effective method to solve problem (P). In this method, we write $h(\cdot)=r\circ\sqrt{(\cdot)}$, so that $s(x_{i})=h(x^{2}_{i})$. Since $r$ is increasing and concave, it follows that $h$ is also concave on $\mathbb{R}_{+}$.
At each iteration, after adding perturbation $\epsilon_{i}^{2}$ to each $h$ and linearizing $h$ at $(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}$, the ${\rm IRL}_{2}$ algorithm solves the following weighted $\ell_{2}$ regularized subproblem (for brevity, we still use $S$ and $T$ for the maps),
$y^{k}=S^{x}(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}u(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2},$
@@
$u(x_{i},\epsilon_{i})=h^{\prime}(x_{i}^{2}+\epsilon_{i}^{2})=\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{2\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}\ \text{ and }\ u(x_{i},0)=\begin{cases}\frac{r^{\prime}(|x_{i}|)}{2|x_{i}|}&\text{ if }x_{i}\neq 0\\ \infty&\text{ if }x_{i}=0.\end{cases}$
@@
(33) $y^{k}_{i}=S_{i}^{x}(x,\epsilon):=\frac{1}{1+\frac{2\lambda}{\beta}u(x_{i}^{k},\epsilon_{i}^{k})}(x_{i}^{k}-\tfrac{1}{\beta}\nabla_{i}f(x^{k})).$
If $u(x_{i}^{k},\epsilon_{i}^{k})=\infty$, we simply set $S_{i}^{x}(x^{k},\epsilon^{k})=0$. The damped iteratively reweighted $\ell_{2}$ algorithm, hereinafter named DIRL2 algorithm, is presented in Algorithm 2.
1:  Input $\beta>\alpha L_{\nabla f}/2$, $\alpha\in(0,1)$, $x_{0}$, $\epsilon_{0}>0$, and $\mu\in(0,1)$.
@@
4:     Reweighting: $u(x_{i}^{k},\epsilon_{i}^{k})=h^{\prime}((x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2})$.
@@
6:     $y^{k}=\arg\min\limits_{y\in\mathbb{R}^{n}}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}u(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2}.$
@@
The fixed point iteration of DIRL2 is then given as $(x^{k+1};\epsilon^{k+1})=T(x^{k},\epsilon^{k})$ with $T=(1-\alpha)I+\alpha S$ and
(34) $T(x,\epsilon):=\begin{bmatrix}(1-\alpha)x+\alpha S^{x}(x,\epsilon)\\ (1-\alpha(1-\mu))\epsilon\end{bmatrix}.$
We have the following result to connect the stationary point of (P) with the fixed point of $T$.
Proposition 4.1.
If $x^{*}$ is a stationary point of (P), then $x^{*}=S^{x}(x^{*},0)$ and $(x^{*};0)=T(x^{*},0)$. Conversely, for any $x^{*}$ satisfying $x^{*}=S^{x}(x^{*},0)$, if $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, then $x^{*}$ is a stationary point of (P).
Proof 4.2.
Suppose $x^{*}$ is a stationary point. It follows from (5) that for $i\in\mathcal{I}^{*}$, $x_{i}^{*}=x_{i}^{*}-\tfrac{1}{\beta}(\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*}))$ implying $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$. Therefore, $x_{i}^{*}=S_{i}^{x}(x^{*},0)$ since $u(x_{i}^{*},0)=\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}$. As for $i\in\mathcal{J}^{*}$, $w(0,0)=\infty$, implying that $0=S_{i}(x^{*},0)$. In summary, $x^{*}=S^{x}(x^{*},0)$.
Now consider $x^{*}=S^{x}(x^{*},0)$. If $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, it suffices to verify (5) is satisfied by $x^{*}$. It follows from $x^{*}_{i}=S_{i}^{x}(x^{*},0),i\in\mathcal{I}^{*}$ that $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$, implying $\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*})=0$, completing the proof.
Remark 4.3.
The converse statement of Proposition 4.1 may not be true when $\mathcal{J}^{*}\neq\emptyset$ and $r^{\prime}(|0^{+}|)<\infty$. Since for any $x^{*}$ that satisfies $x_{i}^{*}=S_{i}^{x}(x^{*},0)$, $i\in\mathcal{I}^{*}$, it is naturally true that $u(0,0)=\infty$ and $S_{i}^{x}(x^{*},0)=0$. However, there is no guarantee that $x^{*}$ satisfies (6). In those cases, $S$ may have two types of fixed points — those are stationary points and others are not. This represents a stark difference from the DIRL1 algorithm, where all fixed points of the iteration map are stationary points.
@@ 4.1 Convergence analysis
Proposition 4.4.
Suppose Assumptions 1-3 hold. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 2 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{\nabla f}}{2}$. It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and the reduction satisfies
$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
Moreover, $\lim\limits_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|=0$ and $\{x^{k}\}\subset\mathcal{L}(F^{0})$.
Proof 4.5.
@@
$r(\sqrt{(x_{i}^{k+1})^{2}+(\epsilon_{i}^{k})^{2}})\leq r(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})+u(x_{i}^{k},\epsilon^{k}_{i})((x_{i}^{k+1})^{2}-(x_{i}^{k})^{2}).$
@@
(35) $\Delta F(x^{k+1},\epsilon^{k+1})\geq\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{\nabla f}}{2}\|x^{k}-x^{k+1}\|^{2}.$
For each subproblem, $y^{k}$ satisfies the optimality condition
(36) $\nabla f(x^{k})+2\lambda U(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})=0.$
where $U(x^{k},\epsilon^{k}):={\rm diag}(u(x_{1}^{k},\epsilon_{1}^{k}),...,u(x_{n}^{k},\epsilon_{n}^{k}))$. Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimality condition (26), the following holds
(37) $\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$
$=$ $\ \nabla f(x^{k})^{T}(x^{k}-x^{k+1})+\lambda\sum\limits_{i=1}^{n}u(x_{i}^{k},\epsilon_{i}^{k})((x_{i}^{k})^{2}-(x_{i}^{k+1})^{2})-\frac{\beta}{2}\|x^{k}-x^{k+1}\|^{2}$
$\geq$ $\ \nabla f(x^{k})^{T}(\alpha x^{k}-\alpha y^{k})+\lambda\sum\limits_{i=1}^{n}u(x_{i}^{k},\epsilon_{i}^{k})(\alpha(x_{i}^{k})^{2}-\alpha(y_{i}^{k})^{2})-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$\geq$ $\ \alpha[\nabla f(x^{k})^{T}(x^{k}-y^{k})+2\lambda\sum\limits_{i=1}^{n}u(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}(x_{i}^{k}-y_{i}^{k})]-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ \alpha[\nabla f(y^{k})+2\lambda U(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})]^{T}(x^{k}-y^{k})+\frac{\alpha\beta(2-\alpha)}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2}.$
Here the first inequality is by $(1-\alpha)(x_{i}^{k})^{2}-\alpha(y_{i}^{k})^{2}\geq(\alpha x_{i}^{k}+(1-\alpha)y_{i}^{k})^{2}$ and the second is by the convexity of $(\cdot)^{2}$. Combining (25) and (37), we get
$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
@@
$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
It follows that $F(x^{k})\leq F(x^{k},\epsilon^{k})\leq F^{0}$, which implies that $\{x^{k}\}\in\mathcal{L}(F^{0})$. By Assumption 3, it is bounded and $\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})>-\infty$. Taking $k\rightarrow\infty$, we have
$(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|=0$.
With the monotonicity of $F$, we derive the global convergence for the ${\rm DIRL}_{2}$ method.
Theorem 4.6.
(Global convergence) Suppose that Assumptions 13 hold, and $\mathcal{J}^{*}=\emptyset$ or $r^{\prime}(|0^{+}|)=\infty$. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{2}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then any $x^{*}\in\Omega$ is a stationary point for (P).
Proof 4.7.
Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. The optimality condition of the $k$th subproblem implies
(38) $-\beta(y_{i}^{k}-x_{i}^{k})=\nabla_{i}f(x^{k})+2\lambda u(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}=\nabla_{i}f(x^{k})+\frac{r^{\prime}(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})}{\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}}}y_{i}^{k}.$
As $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$, $y_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$, $\lim\limits_{k\in\mathcal{S}}\frac{y_{i}^{k}}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}=\text{sign}(x_{i}^{*})$ for $i\in\mathcal{I}^{*}$. Therefore, (5) is satisfied.
If $\mathcal{J}^{*}=\emptyset$, there is nothing to prove. If not, for $i\in\mathcal{J}^{*}$, there is nothing to prove since $|\nabla_{i}f(x^{*})|\leq r^{\prime}(|0^{+}|)$ naturally holds. In general, $x^{*}$ is a stationary point.